Ten consecutive integers are arranged in order from least to greatest. If the sum of the first five integers is 200, what is the sum of the last five integers ?

Answer by checkley77(12844)

X+X+1+X+2+X+3+X+4=200
5X+10=200
5X=200-10
5X=190
X=190/5
X=38 ANS. FOR THE FIRST INTEGER.
38+1=39 ANS. THE SECOND INTEGER.
38+2=40 ANS. FOR THE THIRD INTEGER.
38+3=41 ANS. FOR THE FOURTH INTEGER.
38+4=42 ANS. FOR THE FIFTH INTEGER.
43+44+45+46+47=225 ANS. FOR THE SUM OF THE NEXT FIVE INTEGERS.